OpenStudy (anonymous):

Medal and Fan What is the x-intercept of the line with the equation 1/3x + y = -15 (? , 0) @cj49

OpenStudy (anonymous):

@cj49

OpenStudy (cj49):

-5

OpenStudy (anonymous):

put y = 0 and calculate for x .... that will be the intercept for x

OpenStudy (anonymous):

@cj49 explain your answer

OpenStudy (cj49):

-1/3(15)=-5

OpenStudy (anonymous):

@cj49 the original question is: \[\frac{1}{3}x + y = -15 \\ \text{Put y = 0 :} \\ \frac{1}{3}x = -15\] Now tell me what is the value of x here..??

OpenStudy (anonymous):

Is that your original question @CJBALLER5323 ?? Have I interpreted it right??

OpenStudy (cj49):

-45

OpenStudy (anonymous):

firstly do not give the answers directly and secondly that is too wrong... Please read this http://openstudy.com/code-of-conduct

OpenStudy (cj49):

thank u

OpenStudy (anonymous):

Ur welcome

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@cj49 don't give direct answers, back your solution up with some explanation..

OpenStudy (anonymous):

Like...!! @sushant20902179